An $80,000 recycled paper baling machine was purchased 2 years ago. The projected market value of the machine and the projected net revenue from its operation are shown in the table below. If the company's MARR is 9%, what is the best time for the company to abandon the recycled paper baling project to maximize its present worth?

End of Year: 1 2 3 4

Net Revenue: $20000 $18000 $12000 $3000

Machine Market Value: $35000 $25000 $18000 $10000
Answer #1

Statement showing present worth at the end of different years :

Year Net revenue (a) Market value (b) Present value factor at 9% (c) Present value net revenue (a × c) Accumulated present value of net revenue (d) Present value of market value (e = b × c) Present value of total cash inflows (d + e)
1 20000 35000 0.92 18400 18400 32200 50600
2 18000 25000 0.84 15120 33520 21000 54520
3 12000 18000 0.77 9240 42760 13860 56620
4 3000 10000 0.71 2130 44890 7100 51990

It is beneficial to abandon the machine at the end of year 3 as present worth is maximum in the year 3.

Note : calculations of present value factor

Present value factor = 1/(1+r)n

Where, r = rate of interest and n = number of years.

present value factor at the end of year 1at 9% = 1/(1+0.09)1 = 1/(1.09)1 = 1/(1.09) = 0.92

At the end of year 2 = 1/(1.09)2 = 0.84

At the end of year 3 = 1/(1.09)3 =0.77

At the end of year 4 = 1/(1.09)4 = 0.71